I wasn't sure if this is the correct place to put this, but I need some guidance. I cannot figure out this problem:
Solve the system by substitution.
x+y=4
x-y=2
Thank you for any help you can provide.

Answer by checkley79(3341)

x+y=4
x-y=2
X=Y+2
(Y+2)+Y=4
2Y=4-2
2Y=2
Y=2/2
Y=1 ANS.
X+1=4
X=4-1
X=3 ANS.
PROOF:
3-1=2
2=2